The niagra falls sky wheel has a diameter of 50.5 m and rotates 1 revolution in 2 minutes. The lowest point on the wheel is 2.5m from the ground.

a.)Write an equation that will model the height of a rider over time
b.) use your equation to find the height of a rider after 10 seconds.
Thankyou to anybody who can help me!!!!!!!!

Answer by freakygurrl1231(1)

A.) 30seconds=1/4 of wheel, 1/2=1min, 3/4=1 30 seconds, 2=1 complete rotation.
B.) H=5.3 m


ur welcome. hope i finished in time 4 u!!!! <3